2 1/5 + 3 1/2 + 2 1/2x = 4 3/20
Answer by checkley77(12844)

2 1/5 + 3 1/2 + 2 1/2x = 4 3/20
2.2+3.5+2.5x=4.15
2.5x=4.15-2.2-3.5
2.5x=-1.55
x=-1.55/2.5
x=-.62 ans.